# CAS-ViT: Convolutional Additive Self-attention Vision Transformers for Efficient Mobile Applications

### v1

@@ Methods
Afterward, it passes through four stage encoding layers, with Patch Embedding downsampling 2 times used between each stage, and obtained feature maps of size , and . denotes the feature map channels. Each stage contains stacked blocks as shown in Fig. 3(Lower), and the feature map size remains constant.
The design of block references hybrid networks such as EfficientViT (Liu et al. 2023) and EdgeViT (Pan et al. 2022) and contains three parts with residual shortcuts: Integration subnet, CATM and MLP. Inspired by SwiftFormer (Shaker et al. 2023), Integration sub-network consists of three depth-wise convolution layers activated by ReLU (Glorot, Bordes, and Bengio 2011). We construct a family of lightweight ViT models by varying the number of channels and blocks , and the specific parameter settings refer to Appendix.
Model | Par. | Flops | Throughput(images/s) | Top-1 | Model | Par. | Flops | Throughput(images/s) | Top-1 | ||||
(M) | (M) | GPU | ONNX | ANE | (%) | (M) | (M) | GPU | ONNX | ANE | (%) | ||
SpectFormer-T | 8.88 | 1688 | 2389 | - | - | 76.9 | PoolFormer-S24 | 21.35 | 3394 | 1258 | 36.06 | 3.26 | 80.3 |
MobileOne-S1 | 4.76 | 830 | 4323 | 1366.80 | 20.85 | 77.4 | SpectFormer-XS | 19.61 | 3765 | 1581 | - | - | 80.2 |
FAT-B0 | 4.49 | 710 | 1688 | 50.97 | 6.00 | 77.6 | AFNO-ViT-S/4 | 14.88 | 2840 | 1245 | - | - | 80.9 |
MobileNetv3-L-1.0 | 4.18 | 215 | 6303 | 226.38 | 36.70 | 75.2 | MobileOne-S4 | 14.82 | 2987 | 1468 | 125.20 | 9.70 | 79.4 |
EdgeViT-XXS | 4.07 | 546 | 2438 | 107.16 | 10.31 | 74.4 | PVTv2-B1 | 14.00 | 2034 | 1385 | 21.91 | 3.41 | 78.7 |
EfficientFormerV2-S0 | 3.59 | 400 | 1008 | 73.83 | 19.13 | 75.7 | EfficientViT-M5 | 12.47 | 525 | 5104 | 277.10 | - | 77.1 |
SwiftFormer-XS | 3.47 | 608 | 3184 | 132.69 | 17.11 | 75.7 | SwiftFormer-L1 | 12.05 | 1604 | 2047 | 80.62 | 9.93 | 79.8 |
EMO-2M | 2.38 | 425 | 2672 | 80.11 | 19.07 | 75.1 | GC ViT-XXT | 11.94 | 1939 | 951 | 41.45 | 2.76 | 79.9 |
MobileViT-XS | 2.31 | 706 | 2893 | 129.13 | 17.90 | 75.7 | MobileViTv2-1.5 | 10.56 | 3151 | 1526 | 58.38 | 6.23 | 80.4 |
CAS-ViT-XS | 3.20 | 560 | 3251 | 115.30 | 14.63 | 77.5 | CAS-ViT-M | 12.42 | 1887 | 1566 | 50.39 | 5.49 | 81.4 |
SDT-8-384 | 16.43 | 18802 | - | - | - | 72.3 | ViT-B/16 | 86.42 | 16864 | 705 | 18.27 | 0.93 | 77.9 |
PoolFormer-S12 | 11.90 | 1813 | 2428 | 75.51 | 6.33 | 77.2 | ResNeXt101-64x4d | 83.46 | 15585 | 606 | 36.84 | 2.75 | 79.6 |
MobileOne-S3 | 10.07 | 1902 | 2899 | 652.73 | 16.16 | 80.0 | PoolFormer-S36 | 32.80 | 14620 | 853 | 17.29 | 2.16 | 81.4 |
EfficientViT-M4 | 8.80 | 301 | 6975 | 404.03 | - | 74.3 | SpectFormer-S | 32.03 | 6197 | 966 | - | - | 81.7 |
FAT-B1 | 7.81 | 1186 | 1290 | 37.77 | 4.27 | 80.1 | SDT-8-512 | 29.18 | 33238 | - | - | - | 74.6 |
EdgeViT-XS | 6.78 | 1123 | 2044 | 85.17 | 6.77 | 77.5 | SwiftFormer-L3 * | 28.48 | 4021 | 1309 | 42.23 | 6.4 | 83.0 |
EfficientFormerV2-S1 | 6.18 | 666 | 881 | 55.51 | 14.33 | 79.0 | Swin-T | 28.27 | 4372 | 977 | - | - | 81.3 |
SwiftFormer-S * | 6.09 | 991 | 2607 | 108.98 | 12.90 | 78.5 | PVT-S | 24.10 | 3687 | 933 | 40.62 | 3.08 | 79.8 |
EMO-5M | 5.11 | 883 | 1971 | 53.61 | 10.43 | 78.4 | Twins-SVT-S | 24.06 | 2821 | 1202 | 46.45 | - | 81.7 |
MobileViTv2-1.0 | 4.88 | 1412 | 2429 | 91.40 | 11.52 | 78.1 | Wave-ViT-S | 22.69 | 4391 | 653 | - | - | 82.7 |
CAS-ViT-S | 5.76 | 932 | 2151 | 73.06 | 9.92 | 80.2 | CAS-ViT-T | 21.76 | 3597 | 1084 | 38.67 | 3.83 | 82.3 |
Experiments
ImageNet-1K Classification
Implementation Details
@@
ImageNet-1K (Deng et al. 2009) contains over 1.3 million images across 1,000 natural categories. The dataset covers a wide range of objects and scenes and has become one of the most extensively used datasets with its diversity. We trained the network from scratch without any pre-trained models or additional data. The training strategy follows EdgeNeXt (Maaz et al. 2022), all models are trained on input size of 224224 using AdamW (Loshchilov and Hutter 2018) optimizer for 300 epochs with batch size of 2048. Learning rate is set to 610-3 and the cosine (Loshchilov and Hutter 2016) decay schedule with 20 epochs warmup is used. Label smoothing 0.1 (Szegedy et al. 2016), random resize crop, horizontal flip, RandAugment (Cubuk et al. 2020), multi-scale sampler (Mehta and Rastegari 2021) are enabled, and the momentum of EMA (Polyak and Juditsky 1992) is set to 0.9995 during training. To take full advantage of network effectiveness, we finetuned the model for another 30 epochs at a learning rate of at 384384 resolution with batch size of 64.
@@
We implement the classification model on TIMM 111https://github.com/rwightman/pytorch-image-models based on PyTorch (Paszke et al. 2019) and run on 16 V100 GPUs. In addition, we compile the Torch model into ONNX format and measure the throughput on V100 GPUs and Intel Xeon Gold CPU @ 3.00GHz with batch size 64, respectively. For the mobile side, we compiled it through CoreML library 222https://apple.github.io/coremltools/docs-guides/ and throughput is measured on deployment to iPhone X neural engine.
Backbone | Par.(M) | Flops(G) | mIoU(%) |
CAS-ViT-XS | 6.9 | 24.2 | 36.8 |
PVT-S | 28.2 | 44.5 | 39.8 |
PoolFormer-S24 | 25.1 | 39.3 | 40.3 |
EfficientFormer-L1 | 15.6 | 28.2 | 38.9 |
EdgeViT-XXS | 7.9 | 24.4 | 37.9 |
CAS-ViT-S | 9.4 | 26.1 | 40.3 |
PoolFormer-S36 | 34.6 | 47.5 | 42.0 |
PVTv2-B1 | 17.8 | 34.2 | 42.5 |
EfficientFormerV2-S2 | 16.3 | 27.8 | 42.4 |
SwiftFormer-L1 | 15.5 | 29.7 | 41.4 |
CAS-ViT-M | 15.8 | 31.3 | 43.6 |
PoolFormer-M48 | 77.1 | 82.1 | 42.7 |
PVT-L | 65.1 | 79.6 | 42.1 |
XCiT-S24 P16 | 51.8 | - | 44.6 |
EfficientFormer-L3 | 34.7 | 42.0 | 43.5 |
SwiftFormer-L3 | 31.7 | 42.5 | 43.9 |
CAS-ViT-T | 25.0 | 40.39 | 45.0 |
Results
The experimental outcomes on the ImageNet-1K (Deng et al. 2009) dataset in Table 1 distinctly illustrate the advancements our model introduces in the realm of image classification. Compared to established benchmarks, our approach significantly elevates the precision of classification while adeptly managing the trade-offs between model complexity and computational demand. Notably, the XS and S variants of our model demonstrate an exceptional synergy between size, measured in millions of parameters, and computational efficiency, quantified through Flops, without compromising on the crucial metric of top-1 accuracy. This achievement is pivotal, as it suggests our model’s capability to maintain high accuracy levels even in constrained computational settings, a testament to its optimized architecture. The superior performance of our model, especially in achieving higher top-1 accuracy with fewer parameters and lower Flops compared to traditional models like MobileNetV3 (Howard et al. 2019), underscores its potential for real-world applications where resources are limited. Moreover, the throughput analysis across diverse platforms (GPU, ONNX, and ANE) highlights our model’s adaptability and efficiency, making it an ideal candidate for deployment in various scenarios, from mobile devices to high-end servers, thereby broadening the scope of its applicability in the field of image recognition. For comparisons of extra information like finetuning please refer to the Appendix.
Computational Efficiency
Across all tasks, our models maintain a competitive edge in computational efficiency. This is particularly evident when comparing the Flops and parameter counts, where our models achieve higher or comparable performance metrics with lower computational overheads. For instance, CAS-ViT-M outperforms all listed models in both detection and segmentation tasks while keeping the computational costs competitive, underscoring the effectiveness of our architectural choices.
Object Detection and Instance Segmentation
Implementation Details
@@ Results
Token Mixer | Flops(G) | Top-1(%) | |
1 | +CATM | 1.88 | 80.43 |
2 | CATM Pooling | 1.46 | 78.71 |
3 | CATM W-MSA | 1.96 | 80.41 |
Similarity Function | Flops(G) | Top-1(%) | |
4 | w/o spatial | 1.86 | 79.98 |
5 | w/o channel | 1.88 | 80.22 |
6 | , | 1.87 | 80.04 |
7 | 1.88 | 80.40 |
@@
Taking the pre-trained model on ImageNet-1K as backbone, RetinaNet (Lin et al. 2017) and Mask RCNN (He et al. 2017) are integrated to evaluate the performance of our model on object detection and instance segmentation on MS COCO 2017 (Lin et al. 2014) dataset. Following (Yu et al. 2022), we adopt 1 training strategy and the network is finetuned 12 epochs with learning rate of 210-4 and batch size of 32, and the AdamW (Loshchilov and Hutter 2018) optimizer is used. The training images are resized to 800 pixels on the short side and no more than 1333 pixels on the long side. The models are implemented on MMDeteciton (Chen et al. 2019) codebase.
Results
@@
In the comprehensive evaluation of object detection and instance segmentation on COCO val2017 (Lin et al. 2014) dataset, several backbone models were assessed in terms of their Average Precision (AP) metrics across different scales (small, medium, large) and tasks (detection and segmentation), alongside their computational efficiency indicated by parameters (Par.) and Flops. The analysis primarily focuses on comparing the performance of our proposed models against established models like ResNet-50 (He et al. 2016), PVT-T (Wang et al. 2022), PoolFormer-S12 (Yu et al. 2022), EfficientFormer-L1 (Li et al. 2022b), ResNeXt101-32x4d (Xie et al. 2017), PVTv2-B0 (Wang et al. 2022), PoolFormer-S36 (Yu et al. 2022), and SwiftFormer-L1 (Shaker et al. 2023).
Semantic Segmentation
Implementation Details

### v2

@@ III-D Network Architecture
Afterward, it passes through four stage encoding layers, with Patch Embedding downsampling 2 times used between each stage, and obtained feature maps of size , and . denotes the feature map channels. Each stage contains stacked blocks as shown in Fig. 3(Left Lower), and the feature map size remains constant. We construct a family of lightweight ViT models by varying the number of channels and blocks , and the specific parameter settings refer to Table I.
Model | Blocks | Channels | Params(M) | Flops(M) |
XS | [2, 2, 4, 2] | [48, 56, 112, 220] | 3.20 | 560 |
S | [3, 3, 6, 3] | [48, 64, 128, 256] | 5.76 | 932 |
M | [3, 3, 6, 3] | [64, 96, 192, 384] | 12.42 | 1887 |
T | [3, 3, 6, 3] | [96, 128, 256, 512] | 21.76 | 3597 |
Model | Par. | Flops | Throughput(images/s) | Top-1 | ||
(M) | (M) | GPU | ONNX | ANE | (%) | |
ConvNet-Based | ||||||
ResNeXt101[45] | 83.46 | 15585 | 606 | 36.84 | 2.75 | 79.6 |
MobileNetv3-L-1.0[46] | 4.18 | 215 | 6303 | 226.38 | 36.70 | 75.2 |
MobileOne-S1 [47] | 4.76 | 830 | 4323 | 1366.80 | 20.85 | 77.4 |
MobileOne-S3 [47] | 10.07 | 1902 | 2899 | 652.73 | 16.16 | 80.0 |
MobileOne-S4 [47] | 14.80 | 2978 | 1468 | 125.20 | 9.70 | 81.4 |
MSA-Based | ||||||
EfficientViT-M4 [29] | 8.80 | 301 | 6975 | 404.03 | - | 74.3 |
EfficientViT-M5 [29] | 12.47 | 525 | 5104 | 277.10 | - | 77.1 |
EMO-2M [48] | 2.38 | 425 | 2672 | 80.11 | 19.07 | 75.1 |
EMO-5M [48] | 5.11 | 883 | 1971 | 53.61 | 10.43 | 78.4 |
Swin-T [13] | 28.27 | 4372 | 977 | - | - | 81.3 |
Swin-B [13] | 87.70 | 15169 | 616 | - | - | 83.5 |
EdgeViT-XXS [33] | 4.07 | 546 | 2438 | 107.16 | 10.31 | 74.4 |
EdgeViT-XS [33] | 6.78 | 1123 | 2044 | 85.17 | 6.77 | 77.5 |
ViT-B/16 [24] | 86.42 | 16864 | 705 | 18.27 | 0.93 | 77.9 |
PVT-S [31] | 24.10 | 3687 | 933 | 40.62 | 3.08 | 79.8 |
PVTv2-B1 [17] | 14.00 | 2034 | 1385 | 21.91 | 3.41 | 78.7 |
PVTv2-B3 [17] | 45.2 | 6916 | 561 | 3.48 | 1.18 | 83.2 |
EffiFormerV2-S0 [49] | 3.59 | 400 | 1008 | 73.83 | 19.13 | 75.7 |
EffiFormerV2-S1 [49] | 6.18 | 666 | 881 | 55.51 | 14.33 | 79.0 |
EffiFormerV2-S2 [49] | 12.69 | 1268 | 485 | 31.36 | 7.10 | 81.6 |
GC ViT-XXT [50] | 11.94 | 1939 | 951 | 41.45 | 2.76 | 79.9 |
GC ViT-XT [50] | 19.91 | 2683 | 1072 | 8.01 | 1.75 | 82.0 |
Heterogeneous-MSA-Based | ||||||
PoolFormer-S12 [7] | 11.90 | 1813 | 2428 | 75.51 | 6.33 | 77.2 |
PoolFormer-S24 [7] | 21.35 | 3394 | 1258 | 36.06 | 3.26 | 80.3 |
PoolFormer-S36 [7] | 32.80 | 14620 | 853 | 17.29 | 2.16 | 81.4 |
MobileViT-XS [14] | 2.31 | 706 | 2893 | 129.13 | 17.90 | 75.7 |
MobileViTv2-1.0 [21] | 4.88 | 1412 | 2429 | 91.40 | 11.52 | 78.1 |
MobileViTv2-1.5 [21] | 10.56 | 3151 | 1526 | 58.38 | 6.23 | 80.4 |
MobileViTv2-2.0 [21] | 18.40 | 5577 | 974 | 9.33 | 3.23 | 81.2 |
SwiftFormer-XS [20] | 3.47 | 608 | 3184 | 132.69 | 17.11 | 75.7 |
SwiftFormer-S [20] | 6.09 | 991 | 2607 | 108.98 | 12.90 | 78.5 |
SwiftFormer-L1 [20] | 12.05 | 1604 | 2047 | 80.62 | 9.93 | 79.8 |
SwiftFormer-L3 [20] | 28.48 | 4021 | 1309 | 42.23 | 6.48 | 83.0 |
CAS-ViT | ||||||
CAS-ViT-XS | 3.20 | 560 | 3251 | 115.30 | 14.63 | 78.3/78.7 |
CAS-ViT-S | 5.76 | 932 | 2151 | 73.06 | 9.92 | 80.9/81.1 |
CAS-ViT-M | 12.42 | 1887 | 1566 | 50.39 | 5.49 | 82.8/83.0 |
CAS-ViT-T | 21.76 | 3597 | 1084 | 38.67 | 3.83 | 83.9/84.1 |
Model | Par.(M) | Flops(M) | IN-A | IN-O | IN-R | IN-SK |
PVT-T [31] | 12.85 | 1864 | 7.8 | 18.2 | 33.7 | 21.3 |
PVTv2-B1 [17] | 14.00 | 2034 | 14.7 | 19.3 | 41.8 | 28.9 |
EffiFormer-L1 [54] | 12.28 | 1310 | 10.5 | 22.0 | 42.5 | 30.0 |
PoolFormer-S12 [7] | 11.90 | 1813 | 6.7 | 17.3 | 38.2 | 26.1 |
SwiftFormer-L1 [20] | 12.05 | 1604 | 17.1 | 21.8 | 45.4 | 32.6 |
CAS-ViT-M | 12.42 | 1887 | 22.1 | 23.4 | 45.9 | 33.6 |
Swin-T [13] | 28.27 | 4372 | 21.7 | 21.6 | 41.3 | 29.0 |
PVT-S [31] | 24.10 | 3687 | 17.7 | 20.4 | 40.4 | 27.6 |
PVTv2-B2 [17] | 25.34 | 3886 | 27.9 | 22.3 | 45.1 | 32.8 |
EffiFormer-L3 [54] | 31.39 | 3940 | 23.0 | 24.3 | 46.5 | 34.6 |
EffiFormer-L7 [54] | 82.20 | 10187 | 29.0 | 26.4 | 48.2 | 35.8 |
PoolFormer-S24 [7] | 21.35 | 3394 | 14.2 | 20.0 | 41.7 | 29.9 |
PoolFormer-S36 [7] | 32.80 | 14620 | 18.5 | 20.5 | 42.4 | 31.1 |
SwiftFormer-L3 [20] | 28.48 | 4021 | 26.2 | 24.6 | 48.0 | 35.3 |
CAS-ViT-T | 21.76 | 3597 | 29.6 | 25.4 | 48.5 | 36.0 |
IV Experiments
IV-A Image Classification
IV-A1 Implementation Details
@@
ImageNet-1K [1] contains over 1.3 million images across 1,000 natural categories. The dataset covers a wide range of objects and scenes and has become one of the most extensively used datasets with its diversity. We train the network from scratch without any additional data, and use RegNetY-16GF [55] as teacher model for distillation which was pretrained on ImageNet with 82.9% top-1 accuracy. The training strategy follows EdgeNeXt [41], all models are trained on input size of 224224 using AdamW [56] optimizer for 300 epochs with batch size of 4096. Learning rate is set to 610-3 and the cosine [57] decay schedule with 20 epochs warmup is used. Label smoothing 0.1 [58], random resize crop, horizontal flip, RandAugment [59], multi-scale sampler [14] are enabled, and the momentum of EMA [60] is set to 0.9995 during training. To take full advantage of network effectiveness, we finetuned our model for another 30 epochs at a learning rate of at 384384 resolution with batch size of 64. For specific training hyper-parameters settings, please refer to Table V.
@@
We implement the classification model on TIMM 111https://github.com/rwightman/pytorch-image-models based on PyTorch [61] and run on 16 V100 GPUs. In addition, we compile the Torch model into ONNX format and measure the throughput on V100 GPUs and Intel Xeon Gold CPU @ 3.00GHz with batch size 64, respectively. For the mobile side, we compiled it through CoreML library 222https://apple.github.io/coremltools/docs-guides/ and throughput is measured on deployment to iPhone X neural engine.
@@
dataset | ImageNet-1K | |
variant | CAS-ViT-XS/S/M/T | |
task | 2242 pre-training | 3842 finetuning |
batch size | 4096/4096/4096/2048 | 1024 |
base lr | 6e-3 | 5e-6 |
min lr | 1e-6 | 5e-6 |
lr scheduler | cosine | constant |
training epoch | 300 | 30 |
warmup epoch | 20 | None |
warmup scheduler | linear | None |
optimizer | AdamW | AdamW |
optimizer mome. | ,=0.9,0.999 | ,=0.9,0.999 |
color jitter | 0.4 | 0.4 |
auto-aug | rand-m9-mstd0.5-inc1 | rand-m9-mstd0.5-inc1 |
random-erase prob. | 0/0/0.25/0.25 | 0/0/0.25/0.25 |
random-erase mode | pixel | pixel |
mixup | 0/0/0.8/0.8 | 0/0/0.8/0.8 |
cutmix | 0/0/1.0/1.0 | 0/0/1.0/1.0 |
mixup prob. | 0/0/1.0/1.0 | 0/0/1.0/1.0 |
mixup switch prob. | 0.5 | 0.5 |
drop path rate | 0.0 | 0.0 |
label smoothing | 0.1 | 0.1 |
weight decay | 0.05 | 0.05 |
ema | True | True |
ema decay | 0.9995 | 0.9995 |
teacher model | regnety 160 | regnety 160 |
dist. mode | hard | hard |
dist. alpha | 0.5 | 0.5 |
dist. tau | 1.0 | 1.0 |
IV-A2 Results
The experimental outcomes on the ImageNet-1K [1] dataset in Table II distinctly illustrate the advancements our model introduces in the realm of image classification. Note that for throughput, we report data per frame on mobile device, and report the results on GPUs and ONNX with batch size of 64. Our results are shown in bold for all model variants. The data behind “/” is the result of training 450 epoches on ImageNet-1K [1].
Compared to established benchmarks, our approach significantly elevates the precision of classification while adeptly managing the trade-offs between model complexity and computational demand. Specifically, our CAS-ViT-M/T models achieve top-1 accuracy of 82.8% and 83.9%, with data boosted to 83.0% and 84.1% after further training. Such performance surpasses the recent excellent work on heterogeneous MSA, SwiftFormer [20], by 3.2% and 1.1%, with similar or even fewer parameter numbers. M model outperforms MobileViTv2-2.0 [21] by 1.8% with fewer parameters. In comparison with MSA-based methods is that it outperforms EfficientFormerV2-S2 [49] by 1.4%, as well as our T model outperforms Swin-B [13] by 0.6% with 24% fewer parameters.
Meanwhile, the XS and S variants of our model also demonstrate a superior synergy between the parameter number and the computational efficiency measured by Flops. This achievement demonstrates that our model maintains a high level of performance even under limited computational conditions, which is a testament to its excellent architecture.
IV-A3 Computational Efficiency
Moreover, as shown in Table II, we deployed CAS-ViT to diverse platforms (GPU, ONNX, and ANE) and analyzed its computational efficiency with throughput. The robust model adaptation as well as high efficiency make it an ideal candidate for deployment in various scenarios, from mobile devices to high-end servers, thereby broadening the scope of its applicability in the field of image recognition.
Under similar parameter numbers, our S model has faster throughput and 2.7% higher performance than EMO-5M [48] on all platforms, and our M model has 3.1% higher performance than GCViT-XXT [50]. Under similar throughput, other models tend to require more parameter numbers or lower accuracy, e.g. S model v.s. EdgeViT-XS [33], M model v.s. PVTv2-B1 [17]. Although the throughput falls somewhat short of methods that have a clear advantage in efficient inference, such as EfficientViT [29] and MobileViTv2 [21], CAS-ViT has greater superiority in performance and thus is a bit more favorable in this tradeoff.
In summary, our models maintain a competitive edge in computational efficiency across all platforms. This is particularly evident when comparing the Flops and parameter counts, where our models achieve higher or comparable performance metrics with lower computational overheads.
IV-A4 Effective Robustness
In order to implement a valid assessment in model robustness, we used several additional datasets: ImageNet-A [51] contains examples of misclassification by the ResNet [62], which contains 7500 adversarial-filtered images that can easily lead to model performance degradation. ImageNet-O [51] is used to test the robustness of visual models to out-of-distribution samples, and its evaluation results are reported by Area Under Precision-Recall curve (AUPR). ImageNet-R [52] contains 30000 images from categories such as art, cartoons, etc., and has deductive versions of 200 ImageNet-1K [1] classes. ImageNet-Sketch [53] contains 50000 hand-drawn style images with 50 images of each of the 1000 ImageNet classes.
Our models are not finetuned or processed and are used directly for robustness evaluation and all models are evaluated on an image size of 2242. As shown in Table III, we report comparative results of M and T models, with larger values indicating higher performance. On ImageNet-A our M and T models achieve 22.1% and 29.6%, outperforming other models with similar parameters. On ImageNet-Sketch, our T model has similar accuracy to EfficientFormer-L7 [54] (0.2% higher), while the number of parameters and computational effort is only 26.4%/35.3%.
IV-B Object Detection and Instance Segmentation
IV-B1 Implementation Details
@@
Taking the pre-trained model on ImageNet-1K as backbone, RetinaNet [63] and Mask RCNN [64] are integrated to evaluate the performance of our model on object detection and instance segmentation on MS COCO 2017 [65] dataset. Following [7], we adopt 1 training strategy and the network is finetuned 12 epochs with learning rate of 210-4 and batch size of 32, and the AdamW [56] optimizer is used. The training images are resized to 800 pixels on the short side and no more than 1333 pixels on the long side. The models are implemented on MMDeteciton [66] codebase.
IV-B2 Results
@@
In the comprehensive evaluation on COCO val2017 [65] dataset, several backbone models were assessed in terms of their Average Precision (AP) metrics across different scales (small, medium, large) and tasks (detection and segmentation), alongside their computational efficiency indicated by parameters (Par.) and Flops.
As shown in Table IV, compared to ConvNet, our T model has similar number of parameters as ResNet-50 [62], and we achieve a huge advantage of 5.6 on AP box with RetinaNet as well as 5.5/5.2 on AP box/mask with MaskRCNN. Compared with MSA-based models, our T model outperforms Swin-T with fewer number of parameters, and S model surpasses EfficientFormer-L1 [54] by 1.9/1.3 with MaskRCNN. With a similar parameter number, our T model achieves 43.5/39.6 AP box/mask with MaskRCNN, outperforming the recent excellent work SwiftFormer-L3 [20] by 0.8/0.5. Our model exhibits excellent performance, domain adaptation, and extensibility in object detection and instance segmentation tasks.
IV-C Semantic Segmentation
IV-C1 Implementation Details